OpenStudy (butterflyhope):

Solve: -5y + 2y - 9 + y = -25. Show work please. Thank you! ^-^ :)

OpenStudy (anonymous):

We're not here to do your homework for you- do you not understand this question, or do you want someone to do it for you so you don't have to?

OpenStudy (butterflyhope):

Or you could just help me work it out. I don't really understand how to do it :/

OpenStudy (anonymous):

Also, this question doesn't have an equal sign in it- did you make a typo?

OpenStudy (butterflyhope):

Yeah, it's supposed to be equal xD

OpenStudy (anonymous):

Could you type out the question again?

OpenStudy (butterflyhope):

-5y + 2y - 9 + y = -25

OpenStudy (anonymous):

Alright. You start by combining all the ys into one term.

OpenStudy (anonymous):

So -5y + 2y + y equals...

OpenStudy (cj49):

-5y+2y+y=-2y -2y-9=-25 -2y=-25+9 -2y=-16 y=16/2 y=8

OpenStudy (butterflyhope):

-3y?

OpenStudy (anonymous):

It equals -2y, like cj49 gave you

OpenStudy (butterflyhope):

Ohhhh lol thank you :)

OpenStudy (anonymous):

Since they gave you the answer =_=, I'm sure you understand now

OpenStudy (butterflyhope):

Sorry lol if it makes you feel better I'm not gonna remember it in class xD